7.1 Q4

TITLE Express the confidence interval (0.049,0.117) in the form of p-E<p<p+E. <p< (Type integers or decimals.)
Answer #1

2E = (0.117-0.049) = 0.068

E = 0.034

u = (0.117+0.49)/2 = (0.166)/2 = 0.083

Therefore,

CI: (0.083-0.034 < p < 0.083+0.034)